Question: √ 80 -2 √ 20 + 3 √ 45
Please answer step by step.
Answer in book: 9 √ 5
My answer: 10 √5

Answer by richwmiller(17219)

√ 80 -2 √ 20 + 3 √ 45=
sqrt(80)-2*sqrt(20)+ 3*sqrt(45)=
sqrt(4*4+5)-2*sqrt(4*5)+ 3*sqrt(3*3*5)=
4Sqrt(5)-4sqrt(5)+9 sqrt(5)=
(4-4+9)Sqrt(5)=
9*sqrt(5)